OpenStudy (anonymous):

Katie is finding the standard form of the conic parabola. State her mistake by giving the correct vertex. y=x^2+6x+4 y=x^2+6x+9+4 y=(x+3)^2+4

OpenStudy (anonymous):

When she completed the square, she neglected to balance the equation: \[y = x^2 + 6x + 4 = (x^2 + 6x + 9) + 4 - 9 = (x+3)^2 -5\]

OpenStudy (anonymous):

ohk thnks

OpenStudy (anonymous):

is tht the vertex?

OpenStudy (anonymous):

@eyeh8maf is tht the vertex?

OpenStudy (anonymous):

The vertex of a parabola in standard form of \[y = a(x – h)^2 + k\] is (h,k) so (-3,-5)

OpenStudy (anonymous):

wat?